OpenStudy (anonymous):

Use the function to answer the qustion. f(x) = 5x - 10 what is the value of f-1 (-3) ?

OpenStudy (anonymous):

find the inverse of the function...

OpenStudy (anonymous):

A: 9.4 B: 1.4 C: -5 D: -25

OpenStudy (anonymous):

answer is 7/5

OpenStudy (anonymous):

here y=5x-10 or x=(10+y)/5 hence the reqd answer is (-3+10)/5=7/5=1.4

OpenStudy (anonymous):

@akash123 what?

OpenStudy (anonymous):

answer is 7/5 = 1.4

hartnn (hartnn):

there is an easier way out, just put f(x)= -3 and find x -3 = 5x-10 x=?